# Differentially Private Graph Diffusion with Applications in Personalized PageRanks

### v1

@@ Appendix F Proof of Lemmas
For the above optimization problem, let $\mathbf{x}_{i}=\|\mathbf{h}_{i}\|_{1},\lambda_{1}=\frac{\alpha}{2\alpha-1},\lambda_{2}=\frac{\alpha-1}{\sigma},\lambda_{3}=\frac{\alpha-1}{2\alpha-1}$ and $\lambda_{4}=\frac{\alpha}{\sigma}$. Define $f(\mathbf{x}_{i})=\frac{1}{\alpha-1}\ln(\lambda_{1}\exp(\lambda_{2}\mathbf{x}_{i})+\lambda_{3}\exp(-\lambda_{4}\mathbf{x}_{i})$. Consider the gradient term $\nabla\mathcal{D}_{\alpha}(\mathcal{L}(\bm{0},\sigma)\|\mathcal{L}(\mathbf{h},\sigma))$:
$\frac{\partial\mathcal{D}_{\alpha}(\mathcal{L}(\bm{0},\sigma)\|\mathcal{L}(\mathbf{h},\sigma))}{\partial\mathbf{x}_{i}}=\frac{\partial f(\mathbf{x}_{i})}{\partial\mathbf{x}_{i}}=\frac{\lambda_{1}\lambda_{2}\exp(\lambda_{2}\mathbf{x}_{i})-\lambda_{3}\lambda_{4}\exp(-\lambda_{4}\mathbf{x}_{i})}{\lambda_{1}\exp(\lambda_{2}\mathbf{x}_{i})+\lambda_{3}\exp(-\lambda_{4}\mathbf{x}_{i})}$ (66)
@@
$\frac{\partial^{2}\mathcal{D}_{\alpha}(\mathcal{L}(\bm{0},\sigma)\|\mathcal{L}(\mathbf{h},\sigma))}{\partial\mathbf{x}_{i}^{2}}=\frac{\lambda_{1}\lambda_{3}(\lambda_{2}+\lambda_{4})^{2}\exp(\lambda_{2}\mathbf{x}_{i})\exp(-\lambda_{4}\mathbf{x}_{i})}{\lambda_{1}\exp(\lambda_{2}\mathbf{x}_{i})+\lambda_{3}\exp(-\lambda_{4}\mathbf{x}_{i})}>0,$ (67)
$\frac{\partial^{2}\mathcal{D}_{\alpha}(\mathcal{L}(\bm{0},\sigma)\|\mathcal{L}(\mathbf{h},\sigma))}{\partial\mathbf{x}_{i}\mathbf{x}_{j}}=0,i\neq j.$ (68)
@@
$\mathop{\text{maximize}}\limits_{h}\sum_{i=1}^{|\mathcal{V}|}\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}|\mathbf{h}_{i}|)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma})|\mathbf{h}_{i}|)\;\text{s.t. }\|\mathbf{h}\|_{1}=\rho$ (69)
@@
Next, we use the adjustment method to solve the above objective methods. First, we define $L(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{|\mathcal{V}|})=\sum_{i=1}^{|\mathcal{V}|}f(\mathbf{x}_{i})$ and we fix $\mathbf{x}_{3},\ldots,\mathbf{x}_{|\mathcal{V}|}$. We aim to optimize:
$\mathop{\text{maximize}}\limits_{\mathbf{x}_{1},\mathbf{x}_{2}}L(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\ldots,\mathbf{x}_{|\mathcal{V}|})$ (70)
with respect to $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, considering the constraints $\|\mathbf{h}\|_{1}=\rho$. This is equivalent to
$\mathop{\text{maximize}}\limits_{\mathbf{x}_{1},\mathbf{x}_{2}}\sum_{i=1}^{2}L(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{|\mathcal{V}|}),\;\text{s.t. }\mathbf{x}_{1}+\mathbf{x}_{2}=\rho-\sum_{i=3}^{|\mathcal{V}|}\mathbf{x}_{i}.$ (71)
@@
Define $\mathbf{c}_{t}=\rho-\sum_{i=t}^{\mathcal{V}}\mathbf{x}_{i}$. Since $\mathbf{c}_{3}$ is fixed, and $\mathbf{x}_{2}=\mathbf{c}_{3}-\mathbf{x}_{1}$, the objective function become a univariate function, by calculating the derivative
$\frac{\partial L}{\partial\mathbf{x}_{1}}=\frac{(\lambda_{1}\lambda_{3}\lambda_{4}+\lambda_{1}\lambda_{2}\lambda_{3})(\exp(\lambda_{2}\mathbf{x}_{1}+\lambda_{4}(\mathbf{x}_{1}-\mathbf{c}_{3}))-\exp(\lambda_{2}(\mathbf{c}_{3}-\mathbf{x}_{1})-\lambda_{4}\mathbf{x}_{1}))}{(\lambda_{1}\exp(\lambda_{2}\mathbf{x}_{1}+\lambda_{3}\exp(-\lambda_{4}\mathbf{x}_{1})))(\lambda_{1}\exp(\lambda_{2}(\mathbf{c}_{3}-\mathbf{x}_{1}))+\lambda_{3}\exp(-\lambda_{4}(\mathbf{c}_{3}-\mathbf{x}_{1})))}$ (72)
@@
The derivative $\frac{\partial L}{\partial\mathbf{x}_{1}}$ reaches 0 when $\mathbf{x}_{1}=\frac{\mathbf{c}_{3}}{2}$. Furthermore, $\frac{\partial L}{\partial\mathbf{x}_{1}}<0$ for $\mathbf{x}_{1}<\frac{\mathbf{c}_{3}}{2}$, and $\frac{\partial L}{\partial\mathbf{x}_{1}}>0$ for $\mathbf{x}_{1}>\frac{\mathbf{c}_{3}}{2}$. Thus, the maximum value of the function is attained at the endpoints.
@@
$\mathop{\text{maximize}}\limits_{\mathbf{x}_{1},\mathbf{x}_{2}}L(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\ldots,\mathbf{x}_{|\mathcal{V}|})\leq\max\left[L(\mathbf{c}_{3},0,\mathbf{x}_{3},\ldots,\mathbf{x}_{|\mathcal{V}|}),L(0,\mathbf{c}_{3},\mathbf{x}_{3},\ldots,\mathbf{x}_{|\mathcal{V}|})\right]$ (73)
@@
As the function is symmetric, two endpoints attain the same value, i.e. $L(\mathbf{c}_{3},0,\mathbf{x}_{3},\ldots,\mathbf{x}_{|\mathcal{V}|})=L(0,\mathbf{c}_{3},\mathbf{x}_{3},\ldots,\mathbf{x}_{|\mathcal{V}|})$. Now, we aim to maximize the objective function by each time adjustment two variables and fixed the rest variables unchanged:
@@
$\mathop{\text{maximize}}\limits_{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{|\mathcal{V}|}}L(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\ldots,\mathbf{x}_{|\mathcal{V}|})=\mathop{\text{maximize}}\limits_{\begin{subarray}{c}\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},...,\mathbf{x}_{|\mathcal{V}|}\\ \mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{c}_{3}=\rho\end{subarray}}L(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\ldots,\mathbf{x}_{|\mathcal{V}|})$ (74)
@@
$\leq$ $\mathop{\text{maximize}}\limits_{\mathbf{x}_{3},...,\mathbf{x}_{|\mathcal{V}|}}\mathop{\text{maximize}}\limits_{\begin{subarray}{c}\mathbf{x}_{1},\mathbf{x}_{2}\\ \mathbf{x}_{1}+\mathbf{x}_{2}=\rho-\mathbf{c}_{3}\end{subarray}}L(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\ldots,\mathbf{x}_{|\mathcal{V}|})$ (75)
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\mathop{\text{maximize}}\limits_{\mathbf{x}_{3},\mathbf{x}_{4},...,\mathbf{x}_{|\mathcal{V}|}}L(\rho-\mathbf{c}_{3},0,\mathbf{x}_{3},\ldots,\mathbf{x}_{|\mathcal{V}|})\leq\mathop{\text{maximize}}\limits_{\mathbf{x}_{4},...,\mathbf{x}_{|\mathcal{V}|}}L(\rho-\mathbf{c}_{4},0,0,\ldots,\mathbf{x}_{|\mathcal{V}|})$ (76)
@@
where $(a)$ is from Eq. (73), and for each inequality, we adjust the values of $\mathbf{x}_{1}$ and $\mathbf{x}_{i}$ to maximized the objective function while keep the rest variables fixed. From the above reasoning, we maximized the Rényi divergence over two Laplacians with shift.
F.4 Proof of Lemma B.4.
@@
F.5 Proof of Lemma B.5.
To prove this lemma, we consider tracking the $\infty$-Wasserstein distance of the coupled iterates. Given $\mathscr{D}_{K,\Pi}$ and $\mathscr{D}_{K,\Pi}^{\prime}$,
(1) When $\tau=m$: no randomness is injected for $k$ from $1$ to $\tau$, and from Lemma B.2, we have
$\|\mathbf{s}_{\tau}-\mathbf{s}_{\tau}^{\prime}\|_{1}\leq\frac{(1-\gamma_{\text{max}}^{m})\rho_{\text{diff}}}{(1-\gamma_{\text{max}})}$ (89)
Let $\delta_{\mathbf{s}_{\tau}}$ and $\delta_{\mathbf{s}_{\tau}^{\prime}}$ denote the Dirac measure for two processes at step $\tau$, and therefor,
$W_{\infty}(\delta_{\mathbf{s}_{\tau}},\delta_{\mathbf{s}_{\tau}^{\prime}})\leq\frac{(1-\gamma_{\text{max}}^{m})\rho_{\text{diff}}}{(1-\gamma_{\text{max}})}$ (90)
(2) When $\tau\geq m+1$: recall that for any $k(k\geq m+1)$,
@@
For any step $k(k\geq m+1)$, let $\mu_{k}$ and $\nu_{k}$ denote the distribution of $\mathbf{s}_{k}$ and $\mathbf{s}_{k}^{\prime}$ respectively. Further, define $\tilde{\mu}_{k}$ and $\tilde{\nu}_{k}$ denote the distribution of $\mathbf{S}_{k}=(\xi_{m+1:k},\tilde{\xi}_{m+1:k})$ and $\mathbf{S}_{k}^{\prime}=(\xi_{m+1:k}^{\prime},\tilde{\xi}_{m+1:k}^{\prime})$, respectively. For step $m+1$, we have
$W_{\infty}(\mu_{m+1},\nu_{m+1})=\inf_{\pi_{m+1}\in\Gamma(\mu_{m+1},\nu_{m+1})}\mathop{\mathrm{ess\,sup}}_{(\mathbf{s}_{m+1},\mathbf{s}_{m+1}^{\prime})\sim\pi_{m+1}}\|\mathbf{s}_{m+1}-\mathbf{s}_{m+1}^{\prime}\|_{1}$ (92)
$=$ $\inf_{\tilde{\pi}_{m+1}\in\Gamma(\tilde{\mu}_{m+1},\tilde{\nu}_{m+1})}\mathop{\mathrm{ess\,sup}}_{(\mathbf{S}_{m+1},\mathbf{S}_{m+1}^{\prime})\sim\tilde{\pi}_{m+1}}\|\phi_{m+1}(f(\mathbf{s}_{m}))+\xi_{m+1}^{(1)}+\xi_{m+1}^{(2)}-\phi_{m+1}^{\prime}(f^{\prime}(\mathbf{s}_{m}^{\prime}))-\xi_{m+1}^{\prime(1)}-\xi_{m+1}^{\prime(2)}\|_{1}$ (93)
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\mathop{\mathrm{ess\,sup}}_{(\mathbf{S}_{m+1},\mathbf{S}_{m+1}^{\prime})\sim\tilde{\pi}_{m+1}^{*}}\|\phi_{m+1}(f(\mathbf{s}_{m}))-\phi_{m+1}^{\prime}(f^{\prime}(\mathbf{s}_{m}^{\prime}))\|_{1}\stackrel{{\scriptstyle(b)}}{{\leq}}\gamma_{\text{max}}\|\mathbf{s}_{m}-\mathbf{s}_{m}^{\prime}\|_{1}+\rho_{\text{diff}}$ (94)
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $\frac{(1-\gamma_{\text{max}}^{m})\rho_{\text{diff}}}{(1-\gamma_{\text{max}})}+\rho_{\text{diff}}=\frac{(1-\gamma_{\text{max}}^{m+1})\rho_{\text{diff}}}{1-\gamma_{\text{max}}}:=\lambda_{m+1}$ (95)
where $(a)$ is from selecting a coupling $\tilde{\pi}^{*}$ such that r.v.s $\mathbf{S}_{m+1}$ and $\mathbf{S}_{m+1}^{\prime}$ are identical, $(b)$ is derived from Eq. (53) and $(c)$ is obtained from Lemma B.2.
@@
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\mathop{\mathrm{ess\,sup}}_{(\mathbf{S}_{m+1},\mathbf{S}_{m+1}^{\prime})\sim\tilde{\pi}_{m+1}^{*}}\gamma_{\text{max}}^{\tau-(m+1)}\cdot\|\mathbf{s}_{m+1}-\mathbf{s}_{m+1}^{\prime}\|_{1}+\rho_{\text{diff}}\left(1+\gamma_{\text{max}}+\cdots+\gamma_{\text{max}}^{\tau-m-2}\right)$ (104)
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $\gamma_{\text{max}}^{\tau-(m+1)}\cdot\frac{(1-\gamma_{\text{max}}^{m+1})\cdot\rho_{\text{diff}}}{1-\gamma_{\text{max}}}+\frac{\rho_{\text{diff}}\cdot(1-\gamma_{\text{max}}^{\tau-(m+1)})}{1-\gamma_{\text{max}}}$ (105)

### v4

@@ Appendix F Proof of Lemmas
For the above optimization problem, let $\lambda_{1}=\frac{\alpha}{2\alpha-1},\lambda_{2}=\frac{\alpha-1}{\sigma},\lambda_{3}=\frac{\alpha-1}{2\alpha-1}$ and $\lambda_{4}=\frac{\alpha}{\sigma}$. Define $f(h_{i})=\frac{1}{\alpha-1}\ln(\lambda_{1}\exp(\lambda_{2}\cdot h_{i})+\lambda_{3}\exp(-\lambda_{4}\cdot h_{i})$. Consider the gradient term $\nabla\mathcal{D}_{\alpha}(\mathcal{L}(\bm{0},\sigma)\|\mathcal{L}(\mathbf{h},\sigma))$:
$\frac{\partial\mathcal{D}_{\alpha}(\mathcal{L}(\bm{0},\sigma)\|\mathcal{L}(\mathbf{h},\sigma))}{\partial h_{i}}=\frac{\partial f(h_{i})}{\partial h_{i}}=\frac{\lambda_{1}\lambda_{2}\exp(\lambda_{2}h_{i})-\lambda_{3}\lambda_{4}\exp(-\lambda_{4}h_{i})}{\lambda_{1}\exp(\lambda_{2}h_{i})+\lambda_{3}\exp(-\lambda_{4}h_{i})}$ (52)
@@
$\frac{\partial^{2}\mathcal{D}_{\alpha}(\mathcal{L}(\bm{0},\sigma)\|\mathcal{L}(\mathbf{h},\sigma))}{\partial h_{i}^{2}}=\frac{\lambda_{1}\lambda_{3}(\lambda_{2}+\lambda_{4})^{2}\exp(\lambda_{2}h_{i})\exp(-\lambda_{4}h_{i})}{\lambda_{1}\exp(\lambda_{2}h_{i})+\lambda_{3}\exp(-\lambda_{4}h_{i})}>0,$ (53)
$\frac{\partial^{2}\mathcal{D}_{\alpha}(\mathcal{L}(\bm{0},\sigma)\|\mathcal{L}(\mathbf{h},\sigma))}{\partial h_{i}h_{j}}=0,i\neq j.$ (54)
@@
$\mathop{\text{maximize}}\limits_{\mathbf{h}}\sum_{i=1}^{|\mathcal{V}|}\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}|h_{i}|)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma}|h_{i}|))\;\text{s.t. }\|\mathbf{h}\|_{1}=\rho$ (55)
@@
Next, we use the adjustment method to solve the above objective methods. First, we define $L(h_{1},h_{2},\ldots,h_{|\mathcal{V}|})=\sum_{i=1}^{|\mathcal{V}|}f(h_{i})$ and we fix $h_{3},\ldots,h_{|\mathcal{V}|}$. We aim to optimize:
$\mathop{\text{maximize}}\limits_{h_{1},h_{2}}L(h_{1},h_{2},h_{3},\ldots,h_{|\mathcal{V}|})$ (56)
with respect to $h_{1}$ and $h_{2}$, considering the constraints $\|\mathbf{h}\|_{1}=\rho$. This is equivalent to
$\mathop{\text{maximize}}\limits_{h_{1},h_{2}}\sum_{i=1}^{2}L(h_{1},h_{2},\ldots,h_{|\mathcal{V}|}),\;\text{s.t. }h_{1}+h_{2}=\rho-\sum_{i=3}^{|\mathcal{V}|}h_{i}.$ (57)
@@
Define $c_{t}=\rho-\sum_{i=t}^{\mathcal{V}}h_{i}$. Since $c_{3}$ is fixed, and $h_{2}=c_{3}-h_{1}$, the objective function become a univariate function, by calculating the derivative
$\frac{\partial L}{\partial h_{1}}=\frac{(\lambda_{1}\lambda_{3}\lambda_{4}+\lambda_{1}\lambda_{2}\lambda_{3})(\exp(\lambda_{2}h_{1}+\lambda_{4}(h_{1}-c_{3}))-\exp(\lambda_{2}(c_{3}-h_{1})-\lambda_{4}h_{1}))}{(\lambda_{1}\exp(\lambda_{2}h_{1}+\lambda_{3}\exp(-\lambda_{4}h_{1})))(\lambda_{1}\exp(\lambda_{2}(c_{3}-h_{1}))+\lambda_{3}\exp(-\lambda_{4}(c_{3}-h_{1})))}$ (58)
@@
The derivative $\frac{\partial L}{\partial h_{1}}$ reaches 0 when $h_{1}=\frac{c_{3}}{2}$. Furthermore, $\frac{\partial L}{\partial h_{1}}<0$ for $h_{1}<\frac{c_{3}}{2}$, and $\frac{\partial L}{\partial h_{1}}>0$ for $h_{1}>\frac{c_{3}}{2}$. Thus, the maximum value of the function is attained at the endpoints.
@@
$\mathop{\text{maximize}}\limits_{h_{1},h_{2}}L(h_{1},h_{2},h_{3},\ldots,h_{|\mathcal{V}|})\leq\max\left[L(c_{3},0,h_{3},\ldots,h_{|\mathcal{V}|}),L(0,c_{3},h_{3},\ldots,h_{|\mathcal{V}|})\right]$ (59)
@@
As the function is symmetric, two endpoints attain the same value, i.e. $L(c_{3},0,h_{3},\ldots,h_{|\mathcal{V}|})=L(0,c_{3},h_{3},\ldots,h_{|\mathcal{V}|})$. Now, we aim to maximize the objective function by each time adjustment two variables and fixed the rest variables unchanged:
@@
$\mathop{\text{maximize}}\limits_{h_{1},h_{2},...,h_{|\mathcal{V}|}}L(h_{1},h_{2},h_{3},\ldots,h_{|\mathcal{V}|})=\mathop{\text{maximize}}\limits_{\begin{subarray}{c}h_{1},h_{2},h_{3},...,h_{|\mathcal{V}|}\\ h_{1}+h_{2}+c_{3}=\rho\end{subarray}}L(h_{1},h_{2},h_{3},\ldots,h_{|\mathcal{V}|})$ (60)
@@
$\leq$ $\mathop{\text{maximize}}\limits_{h_{3},...,h_{|\mathcal{V}|}}\mathop{\text{maximize}}\limits_{\begin{subarray}{c}h_{1},h_{2}\\ h_{1}+h_{2}=\rho-c_{3}\end{subarray}}L(h_{1},h_{2},h_{3},\ldots,h_{|\mathcal{V}|})$ (61)
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\mathop{\text{maximize}}\limits_{h_{3},h_{4},...,h_{|\mathcal{V}|}}L(\rho-c_{3},0,h_{3},\ldots,h_{|\mathcal{V}|})\leq\mathop{\text{maximize}}\limits_{h_{4},...,h_{|\mathcal{V}|}}L(\rho-c_{4},0,0,\ldots,h_{|\mathcal{V}|})$ (62)
@@
where $(a)$ is from Eq. (59), and for each inequality, we adjust the values of $h_{1}$ and $h_{i}$ to maximized the objective function while keep the rest variables fixed. From the above reasoning, we maximized the Rényi divergence over two Laplacians with shift.
F.3 Proof of Lemma B.3.
@@
Further, we point out that incorporating the projection operator, i.e., $\varphi_{k}=\phi_{k}\circ f\circ\mathscr{P}_{\mathcal{B}}$, does not affect the contractiveness of the mapping. Indeed, projecting a vector onto the $\ell_{1}$-ball via Euclidean projection yields a nonexpansive operator under the $\ell_{1}$ norm [62], and thus this extension does not alter the bound.
F.4 Proof of Lemma B.4.
To prove this lemma, we consider tracking the $\infty$-Wasserstein distance of the coupled iterates. Given $\mathscr{D}_{K,\Pi}$ and $\mathscr{D}_{K,\Pi}^{\prime}$, for $\tau\geq 1$, recall that for any $k(k\geq 1)$,
@@
For any step $k(k\geq 1)$, let $\mu_{k}$ and $\nu_{k}$ denote the distribution of $\mathbf{s}_{k}$ and $\mathbf{s}_{k}^{\prime}$ respectively. Further, define $\tilde{\mu}_{k}$ and $\tilde{\nu}_{k}$ denote the distribution of $\mathbf{S}_{k}=(\xi_{1:k},\tilde{\xi}_{1:k})$ and $\mathbf{S}_{k}^{\prime}=(\xi_{1:k}^{\prime},\tilde{\xi}_{1:k}^{\prime})$, respectively. For step $1$, we have
$W_{\infty}(\mu_{1},\nu_{1})=\inf_{\pi_{1}\in\Gamma(\mu_{1},\nu_{1})}\mathop{\mathrm{ess\,sup}}_{(\mathbf{s}_{1},\mathbf{s}_{1}^{\prime})\sim\pi_{1}}\|\mathbf{s}_{1}-\mathbf{s}_{1}^{\prime}\|_{1}$ (76)
$=$ $\inf_{\tilde{\pi}_{1}\in\Gamma(\tilde{\mu}_{1},\tilde{\nu}_{1})}\mathop{\mathrm{ess\,sup}}_{(\mathbf{S}_{1},\mathbf{S}_{1}^{\prime})\sim\tilde{\pi}_{1}}\|\phi_{1}(f(\mathbf{s}_{0}))+\xi_{1}^{(1)}+\xi_{1}^{(2)}-\phi_{1}^{\prime}(f^{\prime}(\mathbf{s}_{0}))-\xi_{1}^{\prime(1)}-\xi_{1}^{\prime(2)}\|_{1}$ (77)
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\mathop{\mathrm{ess\,sup}}_{(\mathbf{S}_{1},\mathbf{S}_{1}^{\prime})\sim\tilde{\pi}_{1}^{*}}\|\phi_{1}(f(\mathbf{s}_{0}))-\phi_{1}^{\prime}(f^{\prime}(\mathbf{s}_{0}))\|_{1}\leq\rho_{\text{diff}}$ (78)
where $(a)$ is from selecting a coupling $\tilde{\pi}^{*}$ such that r.v.s $\mathbf{S}_{1}$ and $\mathbf{S}_{1}^{\prime}$ are identical.
@@
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\mathop{\mathrm{ess\,sup}}_{(\mathbf{S}_{1},\mathbf{S}_{1}^{\prime})\sim\tilde{\pi}_{1}^{*}}\gamma_{\text{max}}^{\tau-1}\cdot\|\mathbf{s}_{1}-\mathbf{s}_{1}^{\prime}\|_{1}+\rho_{\text{diff}}\left(1+\gamma_{\text{max}}+\cdots+\gamma_{\text{max}}^{\tau-2}\right)$ (87)
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $\gamma_{\text{max}}^{\tau-1}\cdot\frac{(1-\gamma_{\text{max}})\cdot\rho_{\text{diff}}}{1-\gamma_{\text{max}}}+\frac{\rho_{\text{diff}}\cdot(1-\gamma_{\text{max}}^{\tau-1})}{1-\gamma_{\text{max}}}$ (88)